OpenStudy (anonymous):

Which is the exact value of the expression the square root of 48. − the square root of 75. + the square root of 192.?

OpenStudy (anonymous):

gimmick here is to find the largest perfect square that is a factor of 48

OpenStudy (anonymous):

any ideas for that one ?

OpenStudy (anonymous):

36?

OpenStudy (anonymous):

here is my example for \(75\) the largest perfect square that is a factor of \(75\) is \(25\) because \(75=25\times 3\)

OpenStudy (anonymous):

and therefore \(\sqrt{75}=\sqrt{25\times 3}=\sqrt{25}\sqrt{3}=5\sqrt{3}\)

OpenStudy (anonymous):

no not 36 i did not mean the greatest perfect square less than 48, but rather the greatest perfect square that is a FACTOR of 48

OpenStudy (anonymous):

oh so 12 so it would be \[4\sqrt{3}\]

OpenStudy (anonymous):

not 12 sorry 16

OpenStudy (anonymous):

well yes to \(4\sqrt{3}\)

OpenStudy (anonymous):

figured that is what you meant

OpenStudy (anonymous):

so now we get the idea right? \[\sqrt{48}=4\sqrt{3}\] \[\sqrt{75}=5\sqrt{3}\]

OpenStudy (anonymous):

i am going to bet that \(192\) is some perfect square times \(3\)

OpenStudy (anonymous):

yes so for 192 it would be \[8\sqrt{3}\]

OpenStudy (anonymous):

if \(192=3\times 64\) then yes

OpenStudy (anonymous):

so last job is to combine like terms

OpenStudy (anonymous):

Then the new equation is \[4\sqrt{3}-5\sqrt{3}+8\sqrt{3}\] then combine like terms

OpenStudy (anonymous):

so the answer is \[7\sqrt{3}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thank you so much

OpenStudy (anonymous):

yw